Reminiscences of Arnold Ross

Reflecting on the past I now realize how much I owe to Arnold Ross and what I learned from him, not about mathematics but about how to teach, about setting and keeping standards not only for others but also for oneself. I believe it was in the sixties when I first met him at Notre Dame, and at once I realized his extraordinary gift for motivating people, gently and with understanding. During the years when we had contact I had many occasions to observe his concern for the problems of others, and in particular of those whom he called “youngsters.” He really cared. I find myself lucky to have met in my life such an extraordinary personality.

I have never had the opportunity to participate in one of his summer schools but in discussions with him I learned about the value of such enterprises. Two times he was in Germany and helped us with all his experience and wisdom to establish our “Schülerseminar”, modeled in large part on his summer program. Without his help and his devotion to this task we would never have been able do this here so successfully. Later in the nineties, when I wrote to him about my plans to establish a common seminar with young German and Russian students, he generously donated some funds that helped us realize these plans.

Both my wife Erika and myself remember well Bea Ross who, at the side of her husband, gave him a loving and understanding hand in his devoted work. We are sure that Arnold and Bea will be remembered by the many mathematicians and “youngsters” who had, like us, the privilege to meet the Rosses.

Peter Roquette